OpenStudy (anonymous):

digit problems? system of equations? A two digit number is four times the sum of its digits. The tens digit is 3 less than the units digit. What is the number?

OpenStudy (anonymous):

I have got 2 equations but i don't know if they are right or not... t is the tens digit u is the units (ones) digit 10t+u=4(t+u) t=u-3

OpenStudy (anonymous):

I keep getting 9 but its gotta be a 2 digit number...

OpenStudy (slaaibak):

36

OpenStudy (anonymous):

^ can you show your work please?

OpenStudy (slaaibak):

I guessed, lol. But, I will do it for you.

OpenStudy (anonymous):

oh hahaha okay...

OpenStudy (slaaibak):

10t + u = 4(u + t) t = u - 3 10u - 30 + u = 8u - 12 3u = 18 u = 6 t = 6-3 t=3 therefore it's 36

OpenStudy (anonymous):

thanks so much! you wouldnt mind doing a few others would you? just finding the equations?

OpenStudy (slaaibak):

Sure, I'll help